# Aligning Text, Images, and 3D Structure Token-by-Token

### v1

@@ 3.2.2 What matters in input sequence design?
Figure 4: Hybrid number encodings are robust across training data scales.
Sine-cosine encoding of numbers. While we independently tokenized coordinate values, naively learning embeddings for these tokens may fail to capture the inherent ordering of the numbers (e.g., $2$ is between $1$ and $3$). We investigate whether augmenting the learned embeddings with sine-cosine encodings can enforce numeric ordering relationships. Specifically, we evaluate three encoding strategies: (1) fixed sine-cosine encodings, (2) learned embeddings initialized from scratch, and (3) a hybrid approach where embeddings are learned but augmented with sine-cosine encodings. Fig. 4 plots the effect of these encoding strategies on the recognition task with varying training data sizes. While all methods perform on par in the high data regime, standalone fixed sine-cosine and learned embeddings collapse with low data. Consequently, we adopt the hybrid approach as it demonstrates robustness across data regimes. We show the performance of these encoding strategies across all four tasks in Appendix E.
[…]
@@ 3.2.3 What matters in output sequence design?
Further investigation revealed the issue: the first token. During inference, the model decodes sequentially, with the first token guiding the output. For CLEVR images, this token, representing the top-left corner, was biased toward a few codes – see Fig. 5(a) (blue dash plot) – due to CLEVR’s uniform gray background. This caused overfitting and during inference a wrongly predicted first token caused the decoding to diverge. This finding is not just applicable to CLEVR but to any image set with a more uniform background, often the case in graphic and studio design. Moreover, this trend is even evident with real-world images which we show in Appendix E.
@@
To mitigate this issue, we incorporate a center-token reordering scheme to balance the token distribution at the sequence’s starting position. As shown in Fig. 5(a) (top), the sequence begins with the center token of the image, followed by alternating hops to the left and right until the image is covered. This reordering strategy alleviates the issue since the first token now captures a representative part of the scene, instead of an uninformative background patch. This is shown in Fig. 5(a) and by qualitative examples in Fig. 5(b).
@@
Table 2 shows the impact of center-token reordering and weighted loss for the rendering task. Best performance is achieved when both are combined, also shown qualitatively in Fig. 5(b).
3.3 Generalization to complex object shapes
Above, we validated the effectiveness of the structured 3D modality and developed a cookbook of findings for optimally training unified autoregressive models. We now investigate generalization to complex shapes using ObjaWorld.
@@ 3.3 Generalization to complex object shapes
ObjaWorld features scenes composed of Objaverse assets deitke2023objaverse arranged in diverse layouts. In this section, we focus on two scene categories: park scenes (person, bench, lamppost, bird) and living room scenes (person, sofa, coffee table), featuring realistic textured objects with varied geometry positioned at varying locations and heights. We generate $50,\!000$ park and $50,\!000$ living room scenes to create a training set of $100,\!000$ samples and train rendering and recognition models. Our test set consists of $2,\!000$ park and living room scenes of novel object configurations (locations and poses).
@@
In the recognition task, the model receives a novel scene image and predicts object types, 3D locations, and poses. Kyvo achieves a Jaccard Index of $0.6415$, notably lower than its CLEVR performance ($0.9212$, Table 1), which reflects the greater complexity of ObjaWorld scenes. By comparison, Llama3.2-V, prompted with in-context examples to output the same object attributes, performs near zero – failing to accurately predict $(x,y,z)$ coordinates.
In the rendering task, the model takes a structured list of 3D objects – types, locations, and poses – and generates a corresponding image. This requires linking object types and poses to visual appearance and placing them correctly in 3D space. Fig. 6(a) compares model outputs to ground-truth Blender renderings, showing accurate object types, counts, positions, and poses, though some errors occur in fine pose alignment.
@@
Generalizing to novel scene configurations. ObjaWorld’s training scenes contain scene-specific objects (e.g., bird, bench, lamppost appear only in park scenes). We test on out-of-distribution inputs, such as placing park objects in a living room. Fig. 6(b) shows our model’s ability to generalize to such novel scenarios for the rendering task.
@@
Figure 6: (a) Rendering examples. The model predicts images autoregressively from 3D inputs. Errors include pose mispredictions, e.g., the bird, and missing objects, e.g. the sofa. (b) Novel scene configurations. (c) Chaining tasks. Our recognition model predicts the 3D scene representation for an input image, which is then visualized through both our rendering model and Blender.
@@
Chaining tasks. We show that chaining our recognition (image → 3D) and rendering (3D → image) models enables image reconstruction, providing a test of model robustness. Fig. 6(c) shows two examples: the input image, the reconstruction from our chained models, and a Blender rendering of the predicted 3D scene. The Blender view confirms accurate recognition of object types, poses, and positions, while our reconstructed images closely match the input, with only minor pose errors (e.g., the sofa in the first example).
3.4 Extending the 3D modality for explicit shape representations
In Sections 3.23.3, we demonstrated the effectiveness of our structured 3D modality to align 3D information with images and text. To this end, we focused on type-specific object descriptions (e.g., cube, sphere, cylinder in CLEVR & person, bench, bird etc. in ObjaWorld). We now extend our 3D modality to describe objects with their geometry using 3D shape representations. This requires the model to autoregressively encode and reconstruct geometric details of each object while simultaneously performing scene-level recognition and placement – a substantially more challenging task as it significantly increases the complexity of the prediction and input space.
3.4.1 Vector-quantized 3D shapes
To enable autoregressive models to handle explicit 3D shapes, we require compact sequences encoding geometry and appearance. We adopt Trellis xiang2024structured structured latents (slats), which represent objects as sparse voxels $z=\{(z_{i},p_{i})\}_{i=1}^{L}$, where each $p_{i}$ indexes an active voxel in an $N^{3}$ grid and $z_{i}\in\mathbb{R}^{8}$ encodes local features. Despite their sparsity ($L\ll N^{3}$, typically $L\approx 20\text{k}$ for $N=64$), slats pose challenges for autoregressive modeling: predicting both voxel positions and embeddings, and handling long sequences ($L\approx 20\text{k}$ is intractable during training and decoding)
[…]
[SCENE-START][OBJECT-START][SHAPE]<$v_{1}^{1}$,$v_{2}^{1}$,...,$v_{512}^{1}$>[LOCATION]-0.15 1.05 0.00[POSE]0.00 0.00 3.00[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{2}$,$v_{2}^{2}$,...,$v_{512}^{2}$>[LOCATION]0.25 2.10 $0 0$2.05[OBJECT-END][SCENE-END]
@@ 3.4.1 Vector-quantized 3D shapes
3.4.2 Unified 3D shape and scene understanding
We show that Kyvo, with the shape-augmented 3D modality, can reconstruct and render complex 3D objects and scenes. We build a variant of ObjaWorld with complex objects (e.g., barrel, fire balloon, chicken) from Objaverse deitke2023objaverse , and train recognition and rendering models on 100k image-scene pairs.
The recognition task requires the model to: (1) reconstruct full 3D geometry per object, even under occlusion, and (2) infer each object’s 3D position and pose – all from a single image. Fig. 7(a) shows results on unseen scenes, where our model accurately recovers object geometries and spatial layouts via our structured, object-centric 3D modality. In contrast, Trellis xiang2024structured , a diffusion-based image-to-3D method, trained on Objaverse, that models the scene as a whole, often produces distorted shapes (e.g., deformed bottle) and misaligned layouts (e.g., linearly arranged objects).
In the rendering task, the model inputs the structured 3D modality with embedded shape sequences and outputs the corresponding image. It must map shapes to appearance and position objects accurately by location and pose. Fig. 7(b) compares our model’s outputs with Blender renderings. The model reliably captures object shapes and spatial relationships, with minor distortions in challenging cases (e.g., the occluded cheeseburger behind the basketball in the third image).
Qualitative scaling behavior. Fig. 8 shows that with limited data, the rendering model captures coarse layouts but struggles with object geometry and positioning. As training data increases, object appearance and spatial accuracy improve, yielding more consistent renderings.
Figure 8: Qualitative Scaling Behavior. Training on limited data produces amorphous color blobs lacking semantic coherence. Increasing training data progressively improves geometric precision and spatial relationships.
@@ 3.5 Generalization to real-world recognition
We now evaluate our model’s effectiveness for 3D object recognition in real-world scenes. We conduct experiments on two challenging real-world datasets: Objectron ahmadyan2021objectron , which features indoor and outdoor scenes of diverse object categories (e.g., bicycle, camera, car, cereal box); and ARKitScenes dehghan2021arkitscenes , featuring complex indoor environments of many object categories (e.g., bathtub, bed, cabinet, chair). ARKitScenes presents additional complexity due to its scene density and ground truth annotation noise. Following the train-test splits by brazil2023omni3d , we train a recognition model to detect objects by type and predict their 3D center coordinates and size dimensions in metric space. During training, we augment the data with random horizontal flipping. Table 3 reports the Jaccard Index of our Kyvo and a state-of-the-art 3D object detector Cube R-CNN brazil2023omni3d – we apply a $0.05$ confidence threshold
@@
Table 3: Jaccard-accuracy comparison with Cube R-CNN.
to their predictions, as recommended by the authors. Kyvo significantly outperforms Cube R-CNN with two backbone variants on Objectron and performs on par on the more challenging ARKitScenes dataset. This result demonstrates the potential of a general autoregressive framework that aligns images with the 3D modality.
3.6 Analysis and observations
[…]
@@ 3.6 Analysis and observations
Table 4: Training recipe. From scratch, LoRA vs FFT.
Training recipes. Table 4 compares three approaches for model adaptation: training from scratch, LoRA hu2021lora , and full fine-tuning (FFT). FFT from pre-trained language-only weights yields superior performance even when adapting to image and 3D modalities unseen during pre-training – suggesting effective cross-modal transfer with limited domain-specific data. Notably, LoRA performs worse despite its established efficacy for text-only adaptation, indicating limitations when incorporating entirely new modalities.
[…]
@@
Table 5: Effect of backbone varying size and instruction-tuning.
Instruction-tuned backbones and model sizes. Instruction-tuned backbones outperform or are on par with non-instruction-tuned ones across all tasks, as shown in Table 5. However, increasing the model size from 1B to 3B parameters provides no significant gains, with question-answering performance declining, indicating that the 1B model sufficiently captures our dataset’s complexity while avoiding overfitting.
@@ 4 Conclusion & Limitations
We introduce Kyvo, a autoregressive model that aligns structured 3D with language and vision to support a broad range of 3D visual tasks. Our empirical “cookbook”, based on training 307 models, outlines effective design choices, and we demonstrate how our 3D modality scales to complex shape encodings. We will release code and data.
Limitations: We cover limitations through scaling behavior, qualitative and quantitative results. A key challenge is the limited availability of 3D data. While we show that strong performance and within-domain generalization can be achieved with relatively modest training samples, achieving cross-domain generalization demands larger datasets, which are not readily available in 3D domains. A promising direction is extending Kyvo to handle mixed training data when 3D is not always available as a paired modality. This could enable generalization to new domains, even in the absence of 3D information.
@@ Acknowledgments
We thank Damiano Marsili, Raphi Kang, Ilona Demler, and Ziqi Ma for their valuable feedback. Aadarsh is supported by the Kortschak Scholarship. Georgia is supported by the Powell Foundation, Meta, Google and Amazon.
Appendix Contents
A Dataset details A
A.1 Data generation A.1
A.2 3D scene serialization A.2
A.3 Tokenization A.3
A.4 Task sequence formation A.4
B Additional qualitative examples B
C Evaluation C
C.1 3D scenes C.1
C.2 Images C.2
C.3 Text C.3
D Additional implementation details D
D.1 Image VQGAN architecture D.1
D.2 Encoding of numbers D.2
D.3 3D VQ-VAE training D.3
D.4 Compute resources and time D.4
E Additional experiments and observations E
F Failure cases F
@@ A.1 Data generation
CLEVR. We generate CLEVR scenes using the dataset creation code from johnson2017clevr and render the corresponding images with Blender. Each scene outputs a JSON file describing the scene along with its rendered image.
Rendering. We generate $120,\!000$ unique CLEVR scenes as training data for rendering. The JSON files serve as input for the 3D scene representation, while the corresponding images are used to generate output sequences. Further details are provided in the following sections. For evaluation, we use a test set of $2,\!000$ image-JSON pairs.
Recognition. We use the same training data as in rendering but with reversed input-output roles. Here, images serve as input sequences, while JSON files generate the 3D scene output. For evaluation, we use a test set of $2,\!000$ JSON-image pairs.
Instruction-following. We consider four different types of instructions for 3D scene modification: (1) modifying the appearance of objects, (2) adding new objects, (3) removing objects, and (4) moving an object to a desired location. Using $16$ to $28$ text instruction templates, we generate $20,\!000$ input-output pairs per instruction type that we build on top of the initial $20,\!000$ CLEVR scenes. Specifically, we sample a CLEVR scene, apply an instruction template, and generate the corresponding modified scene. Additionally, for instructions involving object appearance modification, we generate an extra $20,\!000$ pairs that do not reference other objects, ensuring modifications apply solely to uniquely identifiable objects within the scene. This results in a total of $100,\!000$ input-output pairs forming the training set. For evaluation, we sample $500$ input-output pairs per instruction type, creating a test set of $2,\!500$ pairs. This approach enhances dataset diversity, improving the model’s ability to generalize across different instruction types. Example text instructions for each type are listed in Table 6.
Question-answering. For question-answer pair generation, we use the question generation engine by johnson2017clevr that uses functional programs and generate $20,\!000$ question-answer pairs for the training data. For evaluation, we use a test set of $2,\!000$ question-answer pairs.
@@
ObjaWorld. We extend the CLEVR framework to support 3D Blender assets from Objaverse deitke2023objaverse , for inclusion of objects beyond basic geometric shapes such as cubes, cylinders, and spheres. Specifically, we adapt the CLEVR code to include objects like person, bird, bench, etc..
First, for experiments showing generalization to complex object shapes (Section 3.3), we consider two scene setups: park and living room. Park scenes are composed of the assets person, bird, bench, and lamppost, while living room scenes use person, sofa, and table to construct the scenes. For training, we generate $50,\!000$ scenes for each setup, resulting in a total of $100,\!000$ scenes. Our test set comprises $4,\!000$ scenes, evenly split between $2,\!000$ park and $2,\!000$ living room scenes.
For experiments showing extension of the 3D modality for explicit shape representations (Section 3.4), we select 20 complex Objaverse objects like barrel, chicken, cheeseburger, etc., and generate 100,000 training scenes containing 2-3 randomly sampled objects. The test set contains 1,000 unseen scenes.
Objectron. We adhere to the official dataset splits by brazil2023omni3d to construct our training and test sets. For each object in a scene, we extract the category name, 3D center camera coordinates, and object dimensions from the annotation files, generating image–3D scene pairs.
ARKitScenes. Similar to Objectron, we follow brazil2023omni3d and use the provided dataset splits and extract object category labels, 3D center coordinates, and dimensions from annotations to generate image-3D scene pairs.
@@
[SCENE-START][OBJECT-START][SHAPE]<$v_{1}^{1}$,$v_{2}^{1}$,...,$v_{512}^{1}$>[LOCATION]-2.45 0.60 1.20[POSE]0.00 0.00 2.30[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{2}$,$v_{2}^{2}$,...,$v_{512}^{2}$>[LOCATION]2.50 -1.35 0.00[POSE]0.00 0.00 1.55[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{3}$,$v_{2}^{3}$,...,$v_{512}^{3}$>[LOCATION]-1.80 -0.90 0.00[POSE]0.00 0.00 1.45[OBJECT-END][SCENE-END]

### v2

@@ 3.2.2 What matters in input sequence design?
Figure 6: Hybrid number encodings are robust across data scales.
Sine-cosine encoding of numbers. While we independently tokenized coordinate values, naively learning embeddings for these tokens may fail to capture the inherent ordering of the numbers (e.g., $2$ is between $1$ and $3$). We investigate whether augmenting the learned embeddings with sine-cosine encodings can enforce numeric ordering relationships. Specifically, we evaluate three encoding strategies: (1) fixed sine-cosine encodings, (2) learned embeddings initialized from scratch, and (3) a hybrid approach where embeddings are learned but augmented with sine-cosine encodings. Fig.˜6 plots the effect of these encoding strategies on the recognition task with varying training data sizes. While all methods perform on par in the high data regime, standalone fixed sine-cosine and learned embeddings collapse with low data. Consequently, we adopt the hybrid approach as it demonstrates robustness across data regimes. We show the performance of these encoding strategies across all four tasks in the Appendix.
Should the image or 3D come first? We investigate the impact of the ordering between the image and 3D modality for instruction-following and QA tasks. Our experiments reveal that placing the image before the 3D sequence leads to better performance compared to the reverse order. Specifically, we achieve an accuracy of $0.8666$ with the sequence $(I,\text{3D},T_{I})\rightarrow(I,\text{3D})$ compared to $0.8350$ with $(\text{3D},I,T_{I})\rightarrow(\text{3D},I)$. Moreover, we obtain an accuracy of $0.4980$ with $(I,\text{3D},T_{Q})\rightarrow T_{A}$ compared to $0.4720$ with $(\text{3D},I,T_{Q})\rightarrow T_{A}$. This improvement could be attributed to the 3D tokens attending to the entirety of the preceding image tokens, enabling better conditioning and performance.
@@ 3.2.3 What matters in output sequence design?
Further investigation revealed the issue: the first token. During inference, the model decodes sequentially, with the first token guiding the output. For CLEVR images, this token, representing the top-left corner, was biased toward a few codes – see Fig. 7(a) (blue dash plot) – due to CLEVR’s uniform gray background. This caused overfitting and during inference a wrongly predicted first token caused the decoding to diverge. This finding is not just applicable to CLEVR but to any image set with a more uniform background, often the case in graphic design. Moreover, this trend is even evident with real-world images which we show in the Appendix.
@@
To mitigate this issue, we incorporate a center-token reordering scheme to balance the token distribution at the sequence’s starting position, by starting from the center token of the image and alternating hops after that as shown in Fig. 7(a). This reordering means the first token now captures a representative part of the scene, instead of an uninformative background patch. The token distribution and method and results are shown in Fig. 7(a) and Fig. 7(b) respectively.
@@
Tab.˜4 shows the impact of center-token reordering and weighted loss for the rendering task. Best performance is achieved when both are combined, also shown in Fig.˜7(b).
3.3 Generalization to complex scene layouts
Above, we validated our structured 3D modality and developed a cookbook for training unified autoregressive models with tokenized 3D shapes and attributes. We verified our findings for 3D shape encodings on diverse Objaverse objects and for 3D layout attributes (e.g., locations) on CLEVR scenes with known shapes but diverse layouts. We now show these findings generalize to ObjaWorld scenes with complex shapes (from Objaverse assets [deitke2023objaverse]) and diverse layouts.
@@ 3.3 Generalization to complex scene layouts
We focus on two categories: park scenes (person, bench, lamppost, bird) and living room scenes (person, sofa, coffee table), featuring realistic textured objects with varied geometry at diverse locations and heights. We generate 50,00050,\!000 scenes per category (total 100,000100,\!000) to train rendering and recognition models, and 2,0002,\!000 additional scenes with unseen object layouts for testing. The rendering and reconstruction tasks are defined the same as with CLEVR.
@@
For recognition, Kyvo achieves a Jaccard Index of $0.6415$, well below its CLEVR score ($0.9212$, Tab.˜3), reflecting ObjaWorld’s increased complexity. Llama3.2-V, prompted with in-context examples for the same attributes, performs near zero, failing to accurately predict 3D locations.
For rendering, the model renders capture accurate object types, counts, and positions, though some errors occur in fine pose alignment (Fig. 8(a)).
@@
Generalizing to novel scene configurations. We test on out-of-distribution inputs (e.g., park-only objects placed in living room and vice-versa). Fig.˜8(b) shows our model can generalize to such novel scenarios for the rendering task.
@@
Figure 8: (a) Rendering examples. The model predicts images autoregressively from 3D inputs. Errors are largely pose mispredictions, e.g., the bird. (b) Novel scene configurations. (c) Chaining tasks. Our recognition model predicts the 3D scene representation for an input image, which is then visualized through both our rendering model and Blender.
@@
Chaining tasks. Chaining our recognition and rendering models enables image reconstruction to test model robustness. Fig.˜8(c) shows an example of input image, chained-model reconstruction, and Blender rendering of the predicted scene. Blender shows accurate object types, poses, and positions, and the reconstructed images also closely match the input, with only minor pose errors (e.g., the birds).
3.4 Unified 3D shape and scene understanding
Sec.˜3.3 focused on the effectiveness of our structured 3D modality for diverse scene layouts with known complex shapes. Therefore, we used type-specific object descriptions (e.g., “bird” or “person”). Here, we bring together all our findings from Sec.˜3.2, including the learned 3D shape tokenization. We show that Kyvo can reconstruct and render complex 3D objects and scenes, using a variant of ObjaWorld with complex objects (e.g., barrel, chicken) from Objaverse [deitke2023objaverse]. We train recognition and rendering models on 100k image-scene pairs.
@@ 3.4 Unified 3D shape and scene understanding
The recognition model must (1) reconstruct full 3D geometry for each object, and (2) infer each object’s 3D position and pose from a single image. Fig. 9(a) shows results on unseen scenes: our model recovers object geometries and spatial layouts via our structured, object-centric 3D modality, while Trellis [xiang2024structured], a diffusion-based image-to-3D model trained on Objaverse that treats the scene holistically, often yields distorted shapes (e.g., a deformed bottle) and misaligned layouts (e.g., linearly arranged objects).
For rendering, the model takes the structured 3D modality with encoded shape tokens and outputs the corresponding image, mapping shape to appearance and placing objects at the correct positions and poses. Fig. 9(b) compares our outputs with Blender renderings, showing that our model reliably captures shapes and spatial relationships, with only minor distortions in challenging cases (e.g., the occluded cheeseburger behind the basketball in the third image).
Qualitative scaling behavior. Fig. 10 shows that with limited data the rendering model produces amorphous color blobs that lack semantic coherence and captures only coarse layouts, while increasing training data progressively improves object geometry, appearance, and spatial relationships, yielding more accurate and consistent renderings.
Figure 10: Qualitative scaling behavior of rendering model.
@@ 3.5 Generalization to real-world recognition
We now evaluate our model’s effectiveness for 3D object recognition in real-world scenes. We conduct experiments on two challenging real-world datasets: Objectron [ahmadyan2021objectron], which features indoor and outdoor scenes of diverse object categories (e.g., bicycle, camera, car, cereal box); and ARKitScenes [dehghan2021arkitscenes], featuring complex indoor environments of many object categories (e.g., bathtub, bed, cabinet, chair). ARKitScenes presents additional complexity due to its scene density and ground truth annotation noise. Following the train-test splits by [brazil2023omni3d], we train a recognition model to detect objects by type and predict their 3D center coordinates and size dimensions in metric space. During training, we augment the data with random horizontal flipping.
@@
(a) Jaccard Acc. vs Cube R-CNN.
Recipe Rendering()($\downarrow$}) Recognition()($\uparrow$}) Instruction()($\uparrow$}) QA()($\uparrow$})
@@
(b) Training recipe: Scratch vs LoRA vs FFT.
Backbone Rendering()($\downarrow$}) Recognition()($\uparrow$}) Instruction()($\uparrow$}) QA()($\uparrow$})
@@
(c) Effect of backbone size and instruction-tuning.
Table 5: Comparing real-world recognition performance; Quantifying the effect of training recipes and different backbones for Kyvo.
Table 5(a) reports the Jaccard Index of our Kyvo and a state-of-the-art 3D object detector Cube R-CNN [brazil2023omni3d] – we apply a $0.05$ confidence threshold to their predictions, as recommended by the authors. Kyvo significantly outperforms Cube R-CNN with two backbone variants on Objectron and performs on par on the more challenging ARKitScenes dataset. This result demonstrates the potential of a general autoregressive framework that aligns images with the 3D modality.
3.6 Analysis and observations
Training recipes. Table 5(b) compares three approaches for model adaptation: training from scratch, LoRA [hu2021lora], and full fine-tuning (FFT). FFT from pre-trained language-only weights yields superior performance even when adapting to image and 3D modalities unseen during pre-training – suggesting effective cross-modal transfer with limited domain-specific data. Notably, LoRA performs worse despite its established efficacy for text-only adaptation, indicating limitations when incorporating entirely new modalities.
Instruction-tuned backbones and model sizes. Instruction-tuned backbones match or outperform non-instruction-tuned ones across all tasks, as shown in Table 5(c). Increasing model size from 1B to 3B provides no significant gains (even performing notably worse for question-answering), indicating that the 1B model sufficiently captures our dataset’s complexity while avoiding overfitting.
@@ 4 Conclusion & Limitations
We introduce Kyvo, a autoregressive model that aligns structured 3D with language and vision to support a broad range of 3D tasks. Our empirical “cookbook”, based on training 307 models, outlines effective design choices, including for tokenizing 3D scene attributes and complex 3D shapes. We will release code and data.
We cover limitations through scaling behavior, qualitative and quantitative results. A key challenge is the limited availability of 3D data. We show that strong performance and within-domain generalization can be achieved with relatively modest training data, but achieving cross-domain generalization demands larger datasets not readily available. A promising direction is extending Kyvo to handle mixed training data, enabling generalization to new domains even when 3D data is not always available as a paired modality.
@@ Acknowledgments
We thank Damiano Marsili, Raphi Kang, Ilona Demler, and Ziqi Ma for their valuable feedback. Aadarsh is supported by the Kortschak Scholarship. Georgia is supported by the Powell Foundation, Meta through the LLM evaluation research grant, Google, and Amazon.
@@ A.1 Data generation
CLEVR. We generate CLEVR scenes using the dataset creation code from [johnson2017clevr] and render the corresponding images with Blender. Each scene outputs a JSON file describing the scene along with its rendered image.
Rendering. We generate 120,000120,\!000 unique CLEVR scenes as training data for rendering. The JSON files serve as input for the 3D scene representation, while the corresponding images are used to generate output sequences. Further details are provided in the following sections. For evaluation, we use a test set of 2,0002,\!000 image-JSON pairs.
Recognition. We use the same training data as in rendering but with reversed input-output roles. Here, images serve as input sequences, while JSON files generate the 3D scene output. For evaluation, we use a test set of 2,0002,\!000 JSON-image pairs.
Instruction-following. We consider four different types of instructions for 3D scene modification: (1) modifying the appearance of objects, (2) adding new objects, (3) removing objects, and (4) moving an object to a desired location. Using $16$ to $28$ text instruction templates, we generate 20,00020,\!000 input-output pairs per instruction type that we build on top of the initial 20,00020,\!000 CLEVR scenes. Specifically, we sample a CLEVR scene, apply an instruction template, and generate the corresponding modified scene. Additionally, for instructions involving object appearance modification, we generate an extra 20,00020,\!000 pairs that do not reference other objects, ensuring modifications apply solely to uniquely identifiable objects within the scene. This results in a total of 100,000100,\!000 input-output pairs forming the training set. For evaluation, we sample $500$ input-output pairs per instruction type, creating a test set of 2,5002,\!500 pairs. This approach enhances dataset diversity, improving the model’s ability to generalize across different instruction types. Example text instructions for each type are listed in Table 6.
Question-answering. For question-answer pair generation, we use the question generation engine by [johnson2017clevr] that uses functional programs and generate 20,00020,\!000 question-answer pairs for the training data. For evaluation, we use a test set of 2,0002,\!000 question-answer pairs.
@@
ObjaWorld. We extend the CLEVR framework to support 3D Blender assets from Objaverse [deitke2023objaverse], for inclusion of objects beyond basic geometric shapes such as cubes, cylinders, and spheres. Specifically, we adapt the CLEVR code to include objects like person, bird, bench, etc..
First, for experiments showing generalization to complex scene layouts (Section 3.3), we consider two scene setups: park and living room. Park scenes are composed of the assets person, bird, bench, and lamppost, while living room scenes use person, sofa, and table to construct the scenes. For training, we generate 50,00050,\!000 scenes for each setup, resulting in a total of 100,000100,\!000 scenes. Our test set comprises 4,0004,\!000 scenes, evenly split between 2,0002,\!000 park and 2,0002,\!000 living room scenes.
For experiments showing unified 3D shape and scene understanding (Section 3.4), we select 20 complex Objaverse objects like barrel, chicken, cheeseburger, etc., and generate 100,000 training scenes containing 2-3 randomly sampled objects. The test set contains 1,000 unseen scenes.
Objectron. We adhere to the official dataset splits by [brazil2023omni3d] to construct our training and test sets. For each object in a scene, we extract the category name, 3D center camera coordinates, and object dimensions from the annotation files, generating image–3D scene pairs.
ARKitScenes. Similar to Objectron, we follow [brazil2023omni3d] and use the provided dataset splits and extract object category labels, 3D center coordinates, and dimensions from annotations to generate image-3D scene pairs.
Objaverse. For training the 3D VQ-VAE, we use $\sim 168$k Objaverse assets from the Sketchfab subset. However, the computational cost of extracting slats from assets is high using the original TRELLIS pipeline, as it requires extracting 150 renders of the assets which are then passed through the DINOv2 encoder. Instead, we extract a single render of each asset only, and pass it as image-conditioning to the pre-trained TRELLIS slat generator. We use the resultant slats as the inputs to the 3D VQ-VAE encoder during training. We found that the slats generated via this synthetic pipeline are relatively faithful to the original asset. Further, the synthetic slats are effective substitutes in the training data as during evaluation, we use slats extracted from the original TRELLIS pipeline and find that performance transfers well, indicating the distributions of slats are similar with the two methods.
@@
[…]